Bisha Corporation is considering trading a truck with a book value of SAR 52,000 with an estimated five-year life for a new truck that would cost SAR 80,000. The old truck could be sold for SAR 55,000. The new truck has a five-year life with no residual value. The new truck would reduce annual operating costs by SAR 4,300 per year.

Prepare a differential analysis on whether to continue with the old machine (Alternative 1) or purchase the new machine (Alternative 2).
Answer #1

If the new machine is been purchased-

Incremental cost of new machine = 80000 - 55000 = 25000

Annual reduction in operating cost = 4300

Total savings because of new machine = 4300*5 = 21500

There would be a loss of 25000 - 21500 = 3500, if new machine is purchased.

Hence it is better to continue with old machine (Alternative 1).